a+q=11 then q=5,a=?

Grade:12

5 Answers

RAJU
8 years ago
6
KARTHIK
8 years ago
a=6
 
santhosh
8 years ago
6
Nihal Patel
8 years ago
a + q = 11 
q = 5 and a = ? 
Substitue the value of q in the equation: 
a + 5 = 11
a = 11- 5
a = 6
SEELA YUVARAJU
8 years ago
6